ディジタル伝送工学特論 2 単位 (選択)

Advanced Theory of Digital Transmission  高田篤 · 教授/ システム創生工学専攻電気電子創生工学コース電気電子システム講座

【授業目的】実際のディジタル伝送システムの理解に必要な伝送理論を修得し，

簡易な伝送システムの設計手法を理解する．【授業概要】有線伝送システム，無線伝送システム等の実用基本システムの理解

を目的に，符号化�復号化，ディジタル変復調，伝送帯域，雑音と符合誤り等のディジタル伝送理論，及び大容量光ファイバ伝送システム (基幹系，アクセス系) 等の実際の有線/無線ディジタル通信システムの概要を講述する．

【授業形式】講義【キーワード】ディジタル変復調，データ伝送技術，伝送システム，光ファイバ

伝送【先行科目】『通信工学』 (1.0)，『通信応用工学』 (1.0)，『コンピュータネット

ワーク』 (1.0)

【関連科目】『通信工学特論』 (0.5) 【履修要件】学部教育における通信工学を理解していること．【履修上の注意】授業を受ける際には，授業時間と同等の時間の予習と復習を毎回

行うことが，授業の理解と単位取得のために必要である．【到達目標】

1. ディジタル伝送システムの理論と構成を習得し，伝送性能を制限する要因を理解する．

2. 簡易なディジタル伝送システム設計手法を理解する．【授業計画】

1. ディジタル伝送系の概要

2. 情報源符号化と符号化雑音

3. ディジタル多重化技術

4. 伝送路符号

5. 中継伝送と符号誤り

6. ディジタル変復調技術

7. ショット雑音と熱雑音

8. 雑音と符号誤り率

9. 光伝送基本技術

10. 光伝送における雑音

11. 光増幅

12. 波長分割多重伝送

13. WDM 伝送系における性能制限要因 14. 波形等化と符号誤り率

15. 定期試験

16. まとめ

【成績評価基準】レポート 25%, 定期試験 75%．全体で 60%以上で合格．【教科書】プリント【参考書】

⋄ 宮内著「通信システム入門」コロナ社 ⋄ 小林編著「光通信工学 (1)-(2)」コロナ社

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216772 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 高田(E棟C-3, 656-7465, takada@ee) (オフィスアワー: (火)13:30-14:30, ( 木)16:30-17:30)

Advanced Theory of Digital Transmission 2 units (selection)

Atsushi Takada · Professor / Electrical and Electronic Systems, Electrical and Electronic Engineering, Systems Innovation Engineering

Target⟩ Understanding analyzation of digital transmission system and primary

design techniques of a transmission system. Outline⟩ For understanding practical digital wireless/cable transmission systems,

transmission theory on coding, digital modulation/demodulation, equalization, noise, error generation is given. Furthermore, optical fiber transmission systems are discussed. Style Lecture.

Style⟩ Lecture �������⟩ digital modulation, data transmission, transmission system, optical

fiber transmission Fundamental Lecture⟩ “Communication Systems”(1.0), “Applied Communication

Engineering”(1.0), “Computer Networks”(1.0) Relational Lecture⟩ “Advanced Theory of Electrical Communication”(0.5) Requirement⟩ Students are required to have a good understanding of undergraduate-

level communication engineering and related subjects. Notice⟩ have preparation and reviewal for 2 hours on every lecture.

Goal⟩

1. Understanding theory and architecture of digital transmission system and limiting factors of transmission performance.

2. Understanding the techniques designing simple digital transmission system.

Schedule⟩

1. Overview of digital transmission

2. Source coding and coding noise

3. Digital multiplexing

4. Transmission code

5. Repeated transmission and code error

6. Digital modulation/demodulation

7. Shot noise and thermal noise

8. Noise and code error rate

9. Basic technology of optical transmission

10. Noise in optical transmission

11. Optical amplification

12. Wavelength division multiplexing (WDM) transmission

13. Limiting factors of WDM transmission system performance 14. Waveform equalization and error rate

15. Examination

16. Summary

Evaluation Criteria⟩ reports 25%, examination 75%. Total of 60 % is required to

pass the class. Textbook⟩ Prints

Reference⟩

⋄ K. Miyauchi ”Communication System”(Corona Pub.) in Japanese ⋄ I. Kobayashi ed. ”Optical fiber communicaitions (1)(2)”(Corona Pub.) in Japanese

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216772

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ A. Takada(Bldg.E ;C-3, 656-7465, takada@ee) (Office Hour: Tue. 13:30-14:30, Thu.16:30-17:30)